OpenStudy (anonymous):

simplify (2x-9)(x+6)

OpenStudy (anonymous):

do you know how to do this?

OpenStudy (anonymous):

no

OpenStudy (lieutenantgeneral):

-(7x + 54)

OpenStudy (anonymous):

it does not mean "simplify" there is no such mathematical operation as "simplify" it means "multiply" using the distributive property

OpenStudy (anonymous):

\[ (2x-9)(x+6)=2x\times x+6\times 2x-9\times x-9\times 6\]

OpenStudy (anonymous):

then some arithmetic

OpenStudy (kc_kennylau):

\[\Large\begin{array}{cl} &(2x-9)\color{blue}{(x+6)}\\ =&2x\color{blue}{(x+6)}-9\color{blue}{(x+6)}\\ =&2x^2+12x-9x-54\\ =&2x^2+3x-54 \end{array}\]